OpenStudy (anonymous):

give the slope and y-intercept y=-1/5x-6

OpenStudy (anonymous):

y = mx + b m = slope b = intercept

OpenStudy (anonymous):

im a little confused with this one. im not exactly sure how to fill in the equation

OpenStudy (anonymous):

see, m is replaced by -1/5 and b by -6

OpenStudy (anonymous):

i have y=-1/5x+-6 but not sure where to go from there

OpenStudy (anonymous):

see, m is replaced by -1/5 and b by -6 m is slope b is intercept slope = -1/5 intercept = -6

OpenStudy (anonymous):

ok thank you. i was trying to solve the equation instead of just looking at it to find the answer

OpenStudy (anonymous):

ohh okk. you're welcome